OpenStudy (anonymous):

This is important!I need the answer to this! In right triangle ABC, ∠C is a right angle and sin A = sin B. What is m∠A?

OpenStudy (amoodarya):



OpenStudy (anonymous):

It's a multiple choice question. A-30 B-45 C-60 D-75

OpenStudy (amoodarya):

\[\sin A=\frac{a}{c}\\sin B=\frac{b}{c}\\sin A=\sin B \rightarrow \frac{ b }{ c }=\frac{ a }{ c }\\b=a\\so A=B=45^0\]

OpenStudy (anonymous):

Thanks. :D